factor x2+4xy+4y2-16
factor 12(x+1)2-17(x+1)+6

Found 2 solutions by stanbon, AnlytcPhil:
Answer by stanbon(75887)

factor x2+4xy+4y^2-16
= (x+2y)^2-4^2
=(x+2y-4)(x+2y+4)
factor 12(x+1)^2-17(x+1)+6
=[(4(x+1)-3)][(3(x+1)-2]
=[4x+1][3x+1]
Cheers,
Stan H.

Answer by AnlytcPhil(1806)

Factor            x² + 4xy + 4y² - 16

We have to recognize that the first three terms
will factor as a perfect square:
x² + 4xy + 4y² - 16
(x + 2y)(x + 2y) - 16
(x + 2y)² - 16
Now this factors again as the difference of two squares
(x + 2y)² - 4²
[(x + 2y) - 4][(x + 2y) + 4]
[x + 2y - 4][x + 2y + 4]
(x + 2y - 4)(x + 2y + 4)



factor 12(x + 1)² - 17(x + 1) + 6
Replace (x + 1) by a single letter, say W, then
W = (x + 1)
12W² - 17W + 6
(3W - 2)(4W - 3)
Change the parentheses to brackets, then replace the
W's with (x + 1)
[3(x + 1) - 2][4(x + 1) - 3]
[3x + 3 - 2]{4x + 4 - 3]
(3x + 1)(4x + 1)
Edwin
AnlytcPhil@aol.com